OpenStudy (anonymous):

A square pyramid has a slant height of 13 meters and a height of 12 meters. What is the surface area if the square base has sides of length 10 meters? Answer 120 m2 260 m2 320 m2 360 m2

OpenStudy (anonymous):

Surface area of base = 10 * 10 = 100m2 Surface area of four sides = 0.5 * 10 * 13 * 4 = 260 m2 Hence total = 360 m2